\begin{thebibliography}{54}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Albert et~al.(2020)Albert, Penney, Schneier, and
  Siva~Kumar]{Albert2020}
K.~Albert, J.~Penney, B.~Schneier, and R.~S. Siva~Kumar.
\newblock Politics of adversarial machine learning.
\newblock In \emph{Towards Trustworthy ML: Rethinking Security and Privacy for
  ML Workshop, Eighth International Conference on Learning Representations
  (ICLR)}, 2020.

\bibitem[Arora(2019)]{Arora2019}
P.~Arora.
\newblock Decolonizing privacy studies.
\newblock \emph{Television \& New Media}, 20\penalty0 (4):\penalty0 366--378,
  2019.

\bibitem[Brunton and Nissenbaum(2015)]{brunton2015obfuscation}
F.~Brunton and H.~Nissenbaum.
\newblock \emph{Obfuscation: A user's guide for privacy and protest}.
\newblock MIT Press, 2015.

\bibitem[Brunton and Nissenbaum(2017)]{ObfuscationWorkshop2017}
F.~Brunton and H.~Nissenbaum.
\newblock Obfuscation going forward: A research agenda.
\newblock In \emph{Obfuscation Workshop 2017}, 2017.

\bibitem[Buolamwini and Gebru(2018)]{buolamwini2018gender}
J.~Buolamwini and T.~Gebru.
\newblock Gender shades: Intersectional accuracy disparities in commercial
  gender classification.
\newblock In \emph{Conference on fairness, accountability and transparency},
  pages 77--91. PMLR, 2018.

\bibitem[Buolamwini and Raji(2019)]{buolamwini2019actionable}
J.~Buolamwini and I.~D. Raji.
\newblock Actionable auditing: Investigating the impact of publicly naming
  biased performance results of commercial {AI} products.
\newblock Conference on Artificial Intelligence, Ethics, and Society, 2019.

\bibitem[Chandrasekaran et~al.(2021)Chandrasekaran, Gao, Tang, Fawaz, Jha, and
  Banerjee]{Chandrasekaran2021}
V.~Chandrasekaran, C.~Gao, B.~Tang, K.~Fawaz, S.~Jha, and S.~Banerjee.
\newblock Face-off: Adversarial face obfuscation.
\newblock \emph{Proc. Priv. Enhancing Technol.}, 2021\penalty0 (2):\penalty0
  369--390, 2021.

\bibitem[Chen et~al.(2021)Chen, Wang, Zhao, Xue, and Qian]{Chen2021}
L.~Chen, H.~Wang, B.~Z.~H. Zhao, M.~Xue, and H.~Qian.
\newblock Oriole: Thwarting privacy against trustworthy deep learning models.
\newblock \emph{arXiv preprint arXiv:2102.11502}, 2021.

\bibitem[Cherepanova et~al.(2021)Cherepanova, Goldblum, Foley, Duan, Dickerson,
  Taylor, and Goldstein]{CherepanovaGFDD21}
V.~Cherepanova, M.~Goldblum, H.~Foley, S.~Duan, J.~P. Dickerson, G.~Taylor, and
  T.~Goldstein.
\newblock Lowkey: Leveraging adversarial attacks to protect social media users
  from facial recognition.
\newblock In \emph{9th International Conference on Learning Representations,
  {ICLR} 2021, Virtual Event, Austria, May 3-7, 2021}. OpenReview.net, 2021.
\newblock URL \url{https://openreview.net/forum?id=hJmtwocEqzc}.

\bibitem[Das(2020)]{das2020subversive}
S.~Das.
\newblock Subversive {AI}: Resisting automated algorithmic surveillance with
  human-centered adversarial machine learning.
\newblock In \emph{Resistance AI Workshop at NeurIPS 2020}, page~4, 2020.

\bibitem[Eubanks(2018)]{Eubanks2018}
V.~Eubanks.
\newblock \emph{Automating inequality: How high-tech tools profile, police, and
  punish the poor}.
\newblock St. Martin's Press, 2018.

\bibitem[Garvie et~al.(2016)Garvie, Bedoya, and Frankle]{Garvie2016}
C.~Garvie, A.~Bedoya, and J.~Frankle.
\newblock \emph{The perpetual line-up: Unregulated police face recognition in
  America}.
\newblock Georgetown Law, Center on Privacy \& Technology, 2016.

\bibitem[Harvey(2017)]{Harvey2017}
A.~Harvey.
\newblock Hyperface.
\newblock \emph{Adam Harvey, ahprojects. com/hyperface}, 2017.

\bibitem[Hill(2020)]{falsearrest}
K.~Hill.
\newblock Another arrest, and jail time, due to a bad facial recognition match.
\newblock \emph{The New York Times}, 2020.
\newblock URL
  \url{https://www.nytimes.com/2020/12/29/technology/facial-recognition-misidentify-jail.html}.

\bibitem[Howe(2015)]{Howe2015surveillance}
D.~C. Howe.
\newblock Surveillance countermeasures: Expressive privacy via obfuscation.
\newblock \emph{Datafied Research}, 4\penalty0 (1):\penalty0 88--98, 2015.

\bibitem[Johnson(2020)]{ibm2020}
K.~Johnson.
\newblock {IBM walked away from facial recognition. What about Amazon and
  Microsoft?}
\newblock \emph{VentureBeat}, 2020.
\newblock URL
  \url{https://venturebeat.com/2020/06/10/ibm-walked-away-from-facial-recognition-what-about-amazon-and-microsoft/}.

\bibitem[{Katrina Ligett}(2020)]{Ligett2020}
{Katrina Ligett}.
\newblock The {Elephant} in the {Room}: {The} {Problems} that
  {Privacy}-{Preserving} {ML} {Can}’t {Solve}, Dec. 2020.
\newblock URL
  \url{https://slideslive.com/38938421/the-elephant-in-the-room-the-problems-that-privacypreserving-ml-can-t-solve}.

\bibitem[Ko et~al.(2020)Ko, Oleson, Ryan, Register, Xie, Tari, Davidson, Druga,
  and Loksa]{Ko2020}
A.~J. Ko, A.~Oleson, N.~Ryan, Y.~Register, B.~Xie, M.~Tari, M.~Davidson,
  S.~Druga, and D.~Loksa.
\newblock It is time for more critical {CS} education.
\newblock \emph{Communications of the ACM}, 63\penalty0 (11):\penalty0 31--33,
  2020.

\bibitem[Kulynych et~al.(2020)Kulynych, Overdorf, Troncoso, and
  G{\"u}rses]{pots}
B.~Kulynych, R.~Overdorf, C.~Troncoso, and S.~G{\"u}rses.
\newblock Pots: protective optimization technologies.
\newblock In \emph{Proceedings of the 2020 Conference on Fairness,
  Accountability, and Transparency}, pages 177--188, 2020.

\bibitem[Li et~al.(2019)Li, Vincent, Tsai, Kaye, and Hecht]{li2019protest}
H.~Li, N.~Vincent, J.~Tsai, J.~Kaye, and B.~Hecht.
\newblock How do people change their technology use in protest? understanding.
\newblock \emph{Proceedings of the ACM on Human-Computer Interaction},
  3\penalty0 (CSCW):\penalty0 1--22, 2019.

\bibitem[Li and Choi(2021)]{Li2021}
T.~Li and M.~S. Choi.
\newblock Deepblur: A simple and effective method for natural image
  obfuscation.
\newblock \emph{arXiv preprint arXiv:2104.02655}, 2021.

\bibitem[Liu et~al.(2021)Liu, Ding, Shaham, Rahayu, Farokhi, and Lin]{Liu2021}
B.~Liu, M.~Ding, S.~Shaham, W.~Rahayu, F.~Farokhi, and Z.~Lin.
\newblock When machine learning meets privacy: A survey and outlook.
\newblock \emph{ACM Computing Surveys (CSUR)}, 54\penalty0 (2):\penalty0 1--36,
  2021.

\bibitem[{Lujo Bauer}(2021)]{Bauer2021}
{Lujo Bauer}.
\newblock From {Gibbons} to {Fooling} {ML}-based {Face} {Recognition}, 2021.
\newblock URL \url{https://3rd.obfuscationworkshop.org/exhibition/lujo-bauer}.

\bibitem[Lum and Isaac(2016)]{lum2016predict}
K.~Lum and W.~Isaac.
\newblock To predict and serve?
\newblock \emph{Significance}, 13\penalty0 (5):\penalty0 14--19, 2016.

\bibitem[Marwick and boyd(2018)]{marwick2018privacy}
A.~E. Marwick and d.~boyd.
\newblock Understanding privacy at the margins.
\newblock \emph{International Journal of Communication}, 12:\penalty0 9, 2018.

\bibitem[Mayer and Mitchell(2012)]{Mayer2012}
J.~R. Mayer and J.~C. Mitchell.
\newblock Third-party web tracking: Policy and technology.
\newblock In \emph{2012 IEEE Symposium on Security and Privacy}, pages
  413--427. IEEE, 2012.

\bibitem[McDonald and Forte(2020)]{McDonald2020}
N.~McDonald and A.~Forte.
\newblock The politics of privacy theories: Moving from norms to
  vulnerabilities.
\newblock In \emph{Proceedings of the 2020 CHI Conference on Human Factors in
  Computing Systems}, CHI '20, page 1–14, New York, NY, USA, 2020.
  Association for Computing Machinery.
\newblock ISBN 9781450367080.
\newblock \doi{10.1145/3313831.3376167}.
\newblock URL \url{https://doi.org/10.1145/3313831.3376167}.

\bibitem[Mirshghallah et~al.(2020)Mirshghallah, Taram, Vepakomma, Singh,
  Raskar, and Esmaeilzadeh]{Mirshghallah2020}
F.~Mirshghallah, M.~Taram, P.~Vepakomma, A.~Singh, R.~Raskar, and
  H.~Esmaeilzadeh.
\newblock Privacy in deep learning: A survey.
\newblock \emph{arXiv preprint arXiv:2004.12254}, 2020.

\bibitem[Moravec(2019)]{racistmath}
E.~R. Moravec.
\newblock Do algorithms have a place in policing?
\newblock \emph{The Atlantic}, 2019.
\newblock URL
  \url{https://www.theatlantic.com/politics/archive/2019/09/do-algorithms-have-place-policing/596851/}.

\bibitem[Muller and Druin(2012)]{Muller2012}
M.~J. Muller and A.~Druin.
\newblock Participatory design: the third space in human--computer interaction.
\newblock \emph{The Human--Computer Interaction Handbook}, pages 1125--1153,
  2012.

\bibitem[Nanayakkara et~al.(2021)Nanayakkara, Hullman, and
  Diakopoulos]{BroaderImpact}
P.~Nanayakkara, J.~Hullman, and N.~Diakopoulos.
\newblock Unpacking the expressed consequences of {AI} research in broader
  impact statements.
\newblock In \emph{Proceedings of the 2021 AAAI/ACM Conference on AI, Ethics,
  and Society}, AIES '21, page 795–806, New York, NY, USA, 2021. Association
  for Computing Machinery.
\newblock ISBN 9781450384735.
\newblock \doi{10.1145/3461702.3462608}.
\newblock URL \url{https://doi.org/10.1145/3461702.3462608}.

\bibitem[Nissenbaum(2009)]{Nissenbaum2009}
H.~Nissenbaum.
\newblock \emph{Privacy in Context: Technology, Policy, and the Integrity of
  Social Life}.
\newblock Stanford University Press, USA, 2009.
\newblock ISBN 0804752370.

\bibitem[Nissenbaum and Daniel(2009)]{nissenbaum2009trackmenot}
H.~Nissenbaum and H.~Daniel.
\newblock Trackmenot: Resisting surveillance in web search.
\newblock 2009.

\bibitem[Papernot et~al.(2018)Papernot, McDaniel, Sinha, and
  Wellman]{Papernot2018}
N.~Papernot, P.~McDaniel, A.~Sinha, and M.~P. Wellman.
\newblock Sok: Security and privacy in machine learning.
\newblock In \emph{2018 IEEE European Symposium on Security and Privacy
  (EuroS\&P)}, pages 399--414. IEEE, 2018.

\bibitem[Polack(2020)]{polack2020beyond}
P.~Polack.
\newblock Beyond algorithmic reformism: Forward engineering the designs of
  algorithmic systems.
\newblock \emph{Big Data \& Society}, 7\penalty0 (1):\penalty0
  2053951720913064, 2020.

\bibitem[Radiya-Dixit and Tram{\`e}r(2021)]{Radiya2021}
E.~Radiya-Dixit and F.~Tram{\`e}r.
\newblock Data poisoning won't save you from facial recognition.
\newblock \emph{arXiv preprint arXiv:2106.14851}, 2021.

\bibitem[Rajabi et~al.(2021)Rajabi, Bobba, Rosulek, Wright, and
  Feng]{Rajabi2021}
A.~Rajabi, R.~B. Bobba, M.~Rosulek, C.~Wright, and W.-C. Feng.
\newblock On the (im) practicality of adversarial perturbation for image
  privacy.
\newblock \emph{Proceedings on Privacy Enhancing Technologies}, 2021.

\bibitem[Raji et~al.(2020)Raji, Gebru, Mitchell, Buolamwini, Lee, and
  Denton]{raji2020}
I.~D. Raji, T.~Gebru, M.~Mitchell, J.~Buolamwini, J.~Lee, and E.~Denton.
\newblock Saving face: Investigating the ethical concerns of facial recognition
  auditing.
\newblock In \emph{Proceedings of the AAAI/ACM Conference on AI, Ethics, and
  Society}, pages 145--151, 2020.

\bibitem[Roberts(2019)]{Roberts2019digitizing}
D.~E. Roberts.
\newblock Digitizing the carceral state.
\newblock 2019.

\bibitem[Rose(2021)]{Rose2021}
K.~Rose.
\newblock Paint {Box} vs {Black} {Box}: {Using} {Art} to {Bring}
  {Accountability} to {AI} and {Surveillance}, Mar. 2021.

\bibitem[Sharif et~al.(2016)Sharif, Bhagavatula, Bauer, and Reiter]{Sharif2016}
M.~Sharif, S.~Bhagavatula, L.~Bauer, and M.~K. Reiter.
\newblock Accessorize to a crime: Real and stealthy attacks on state-of-the-art
  face recognition.
\newblock In \emph{Proceedings of the 2016 ACM SIGSAC Conference on Computer
  and Communications Security}, pages 1528--1540, 2016.

\bibitem[Shen et~al.(2019)Shen, Zhu, and Ma]{Shen2019}
J.~Shen, X.~Zhu, and D.~Ma.
\newblock Tensorclog: An imperceptible poisoning attack on deep neural network
  applications.
\newblock \emph{IEEE Access}, 7:\penalty0 41498--41506, 2019.

\bibitem[Sloan et~al.(2020)Sloan, Moss, Awomolo, and
  Forlano]{Sloan2020participation}
M.~Sloan, E.~Moss, O.~Awomolo, and L.~Forlano.
\newblock Participation is not a design fix for machine learning (pp. 1--7).
\newblock In \emph{Proceedings of the International Conference on Machine
  Learning, Vienna, Austria}, 2020.

\bibitem[Solove(2008)]{Solove2008}
D.~J. Solove.
\newblock Understanding privacy.
\newblock 2008.

\bibitem[Stark(2019)]{Stark2019}
L.~Stark.
\newblock Facial recognition is the plutonium of {AI}.
\newblock \emph{XRDS: Crossroads, The ACM Magazine for Students}, 25\penalty0
  (3):\penalty0 50--55, 2019.

\bibitem[{Stop LAPD Spying Coalition}(2018)]{stop2014spying}
{Stop LAPD Spying Coalition}.
\newblock {Before the bullet hits the body: Dismantling predictive policing in
  Los Angeles}, 2018.

\bibitem[Stroud(2021)]{Stroud2021}
M.~Stroud.
\newblock Heat listed.
\newblock \emph{The Verge}, 2021.
\newblock URL \url{https://www.theverge.com/22444020/heat-listed-csk-entry}.

\bibitem[Tapper(2020)]{dazzleclub}
J.~Tapper.
\newblock Hiding in plain sight: activists don camouflage to beat met
  surveillance.
\newblock \emph{The Guardian}, 2020.
\newblock URL
  \url{https://www.theguardian.com/world/2020/feb/01/privacy-campaigners-dazzle-camouflage-met-police-surveillance}.

\bibitem[Thys et~al.(2019)Thys, Van~Ranst, and Goedeme]{Thys2019}
S.~Thys, W.~Van~Ranst, and T.~Goedeme.
\newblock Fooling automated surveillance cameras: Adversarial patches to attack
  person detection.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition (CVPR) Workshops}, June 2019.

\bibitem[Vincent et~al.(2019)Vincent, Hecht, and Sen]{vincent2019data}
N.~Vincent, B.~Hecht, and S.~Sen.
\newblock “{Data Strikes}”: Evaluating the effectiveness of a new form of
  collective action against technology companies.
\newblock In \emph{The World Wide Web Conference}, pages 1931--1943, 2019.

\bibitem[Vincent et~al.(2021)Vincent, Li, Tilly, Chancellor, and
  Hecht]{vincent2021leverage}
N.~Vincent, H.~Li, N.~Tilly, S.~Chancellor, and B.~Hecht.
\newblock Data leverage: A framework for empowering the public in its
  relationship with technology companies.
\newblock In \emph{Proceedings of the 2021 ACM Conference on Fairness,
  Accountability, and Transparency}, pages 215--227, 2021.

\bibitem[Whitney et~al.(2021)Whitney, Naval, Quepons, Singh, Rick, and
  Irani]{whitney2021hci}
C.~D. Whitney, T.~Naval, E.~Quepons, S.~Singh, S.~R. Rick, and L.~Irani.
\newblock {HCI} tactics for politics from below: Meeting the challenges of
  smart cities.
\newblock In \emph{Proceedings of the 2021 CHI Conference on Human Factors in
  Computing Systems}, pages 1--15, 2021.

\bibitem[Zong(2020)]{Zong}
J.~Zong.
\newblock From individual consent to collective refusal: Changing attitudes
  toward (mis)use of personal data.
\newblock \emph{XRDS}, 27\penalty0 (2):\penalty0 26–29, Dec. 2020.
\newblock ISSN 1528-4972.
\newblock \doi{10.1145/3433140}.
\newblock URL \url{https://doi.org/10.1145/3433140}.

\bibitem[Zuboff(2018)]{Zuboff2018}
S.~Zuboff.
\newblock \emph{The Age of Surveillance Capitalism: The Fight for a Human
  Future at the New Frontier of Power}.
\newblock 1st edition, 2018.
\newblock ISBN 1610395697.

\end{thebibliography}

\documentclass[letterpaper]{article}

\pdfoutput=1

\usepackage[preprint]{neurips_2021}
\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage[hyperindex,breaklinks]{hyperref}
\usepackage{url}   
\usepackage{microtype}      

\setcitestyle{authoryear,open={(},close={)}}

\title{Addressing Privacy Threats from Machine Learning}

\author{%
  Mary Anne Smart \\
  Department of Computer Science \& Engineering\\
  University of California San Diego\\
  \texttt{msmart@ucsd.edu} \\
}

\begin{document}

\maketitle

\begin{abstract}
  Every year at NeurIPS, machine learning researchers gather and discuss exciting applications of machine learning in areas such as public health, disaster response, climate change, education, and more\footnote{These topics are all the subjects of \href{https://blog.neurips.cc/2021/08/20/neurips-2021-workshop-announcement/}{NeurIPS 2021 workshops}.}. However, many of these same researchers are expressing growing concern about applications of machine learning for surveillance \citep{BroaderImpact}. This paper presents a brief overview of strategies for resisting these surveillance technologies and calls for greater collaboration between machine learning and human-computer interaction researchers to address the threats that these technologies pose.
\end{abstract}

\section{Introduction}

A rapidly growing research area in computer science is that of privacy in machine learning \citep{Papernot2018, Mirshghallah2020, Liu2021}. Differentially-private machine learning algorithms and other privacy-preserving methods offer the opportunity to learn from sensitive datasets while limiting what can be revealed about any particular individual. However, some machine learning systems violate privacy---not as some unintended side-effect, but by design. The tools from privacy-preserving machine learning are little use against machine learning models that are designed for biometric or behavioral profiling \citep{Ligett2020}; different approaches are needed.

Although a variety of definitions, frameworks, and taxonomies have been proposed, there is no single, universally agreed-upon definition of privacy \citep{Solove2008, Nissenbaum2009, Arora2019, McDonald2020}. This paper focuses on the specific set of privacy-related harms that are perpetrated or exacerbated by machine learning systems. In an age where powerful algorithms can process large datasets at high speeds, it becomes harder to find privacy by hiding in a crowd. Facial recognition algorithms can identify people in public spaces \citep{Garvie2016, Stark2019}; targeted advertising algorithms can exploit detailed user profiles to shape behavior at scale \citep{Zuboff2018}; and predictive policing algorithms can single out individuals for targeted surveillance \citep{Stroud2021}. These are the kinds of novel threats to privacy that machine learning systems enable. 

A number of strategies have emerged for resisting machine learning systems that threaten privacy. This work provides an overview of such strategies and argues that addressing these privacy challenges effectively will require greater collaboration between the NeurIPS community and experts in the field of human-computer interaction (HCI). This paper discusses two overarching approaches for addressing threats to privacy posed by machine learning (although in some cases the line between the two may be blurred). The first approach involves challenging the data that feeds the malicious machine learning model, either through some form of obfuscation or through withholding information. The second approach involves challenging the malicious model directly, perhaps by applying pressure to force the retirement of the model or by banning the harmful technology altogether. With either approach, there is a role for computer scientists to play.

\section{Challenging Data}
Machine learning systems run on data. First, data is required to train a machine learning model. Next, after the model has been deployed, new data is fed into the model to produce predictions. In practice, these stages of training and deployment may take place iteratively; for example, the deployed model may periodically be retired and replaced by a new model retrained on the latest data. One way to resist a machine learning system is to interfere with the data that feeds it. There are two main strategies that fall under this umbrella: obfuscation and withholding.

\subsection{Obfuscation}

One way to evade machine learning surveillance systems is by strategically altering either 1) the data used to make predictions or 2) the data used to train the system. As an example of the first strategy, \cite{Sharif2016} designed glasses that could fool facial recognition systems. The researchers exploited the fact that machine learning models tend to be vulnerable to \textit{adversarial examples}. An adversarial example is a datapoint that has been modified slightly--such that a human would not perceive any difference but that a machine learning model misclassifies it. In recent years, researchers have developed a number of strategies for evading facial recognition using adversarial examples \citep{Harvey2017, Thys2019, CherepanovaGFDD21, Li2021, Chandrasekaran2021}. These approaches aim to help individuals avoid surveillance but typically lack strong guarantees \citep{Rajabi2021, Bauer2021, Chen2021, Radiya2021}.

An alternative to altering data at the time of prediction is to alter the data used to train the underlying machine learning model. This style of attack is known as a \textit{data poisoning attack}. For example, TensorClog produces altered images that are designed to reduced the accuracy of deep learning models \citep{Shen2019}. Another example of data poisoning for anti-surveillance purposes can be found at Adversarial Fashion\footnote{\url{https://adversarialfashion.com/}}, a shop that sells clothing designed to trigger automated license plate readers and "[inject] junk data" \citep{Rose2021}. 

Although the above examples focus on image classification tasks, similar obfuscation tactics have been used for resisting web tracking, loyalty card-based tracking, and more \citep{brunton2015obfuscation, pots}. While obfuscation can be of practical utility, it can also have an \textit{expressive} function \citep{nissenbaum2009trackmenot, Howe2015surveillance}. For example, members of the "Dazzle Club" paint their faces with unusual makeup and conduct walks through the streets of London \citep{dazzleclub}. Although their makeup is designed to trick facial recognition systems, the group's real purpose is to protest police use of facial recognition. 

Adversarial examples and data poisoning attacks are ongoing topics of study within the machine learning community. If these technologies are to be adopted as anti-surveillance tools, many practical and ethical challenges remain \citep{brunton2015obfuscation, ObfuscationWorkshop2017, das2020subversive, Albert2020}. How can such tools be made more accessible? How should such tools be evaluated? How should the risks of using obfuscation be communicated? Addressing these questions will require collaborations among machine learning experts, HCI researchers, activists, and other stakeholders. 

\subsection{Withholding Data} 
An alternative to altering the data used by machine learning surveillance systems is to withhold the data. One simple example of this strategy is the use of privacy-enhancing technologies that block web tracking \citep{Mayer2012}. Withholding data by using tracker-blocking browser extensions may provide some privacy protection for individuals. However, data can also be withheld at the collective level \citep{vincent2021leverage}. For example, a \textit{data strike} is a kind of digital boycott that can be used to apply public pressure to technology companies \citep{vincent2019data}. A related way of withholding data is through \textit{protest non-use}; many people have simply stopped using certain platforms due to privacy and surveillance-related concerns \citep{li2019protest}. Dazzle club walks, data strikes, and protest non-use all go one step beyond strategies that obfuscate or withhold data solely as a way of evading surveillance; rather, they use these data-focused strategies as starting points to launch broader critiques and campaigns against the surveillance systems themselves.

\section{Challenging Models}
Although the data-oriented approaches described so far offer some help in resisting machine learning surveillance systems, in many cases, policy approaches offer a more satisfactory solution. For example, while a variety of strategies exist for evading or fooling facial recognition, banning particular uses of facial recognition would make these strategies unnecessary. There exist a wide range of forms that regulation can take and a wide range of roles that computer scientists may play in the process.

One technique that can apply pressure to companies that develop and sell surveillance technologies is \emph{auditing}. In the case of facial recognition, audits by researchers have helped call attention to the harm that this technology can cause. It is now well-established that many facial recognition systems perform poorly on darker-skinned subjects \citep{buolamwini2018gender, buolamwini2019actionable}; these accuracy issues have led to multiple wrongful arrests of Black men in the United States \citep{falsearrest}. The audits that uncovered racial bias in facial recognition systems placed pressure on companies that sell facial recognition technology to police. In fact, in June 2020, several major companies stopped selling facial recognition technology due to mounting pressure from researchers, protestors, and their own employees \citep{ibm2020}. Nevertheless, audits have limitations; for example, audits may inadvertently "normalize tasks that are inherently harmful to certain communities" \citep{raji2020}.

Some proprietary technologies cannot be easily audited because access to the technology is restricted.  Nevertheless, it is sometimes possible to reverse engineer a system and show that it can produce societal harm. For example, \cite{lum2016predict} simulate the effects of predictive policing systems and show how predictive policing can magnify existing biases and lead to the over-policing of low-income communities and communities of color. Many of the same risks involved with traditional audits are at play here as well. For example, a narrow focus on particular metrics could simply result in shifting goal posts -- what \cite{polack2020beyond} refers to as \textit{algorithmic reformism}. Therefore, when conducting algorithmic audits or when reverse engineering particular algorithms, it is important to consider the scope of the critique one hopes to make.

In some cases, researchers have partnered with community organizations in order to push back against surveillance technologies. Sometimes mathematical or technical language is used to insulate surveillance technologies from public criticism \citep{stop2014spying}. For example, an LAPD spokesperson infamously said of predictive policing: "It is math, not magic, and it is not racist" \citep{racistmath}. Technologists can debunk the myth that critics of surveillance technologies simply don't understand the math; they can also help parse technical language and demystify machine learning algorithms. However, it is important that researchers approach partnerships with humility; community organizers will bring their own areas of expertise to the table \citep{whitney2021hci}.

Finally, it is important to acknowledge the academic community's complicity in building and upholding the very forms of surveillance discussed in this paper \citep{stop2014spying}. \cite{Ko2020} argue that computer science educators have a responsibility to make the role of computing in injustice visible. At some universities, students have taken it upon themselves to educate each other about the harmful role of computing within US Immigration and Customs Enforcement\footnote{\url{https://notechforice.com}} \citep{Zong}. These kinds of efforts will help the next generation of computer scientists think critically about the consequences of the technologies they create.

\section{Conclusion}
This paper has outlined a range of strategies for resisting surveillance technologies powered by machine learning. It concludes by echoing \cite{das2020subversive} in calling for participatory methods when designing anti-surveillance technologies such as those discussed in this paper. While participatory methods are relatively common in HCI research \citep{Muller2012}, they have received less attention from the machine learning community\footnote{Some exceptions include the \href{https://participatoryml.github.io/}{Participatory Approaches to Machine Learning ICML 2020 Workshop} and the \href{https://sites.google.com/view/resistance-ai-neurips-20/home}{Resistance AI NeurIPS 2020 Workshop}.}. The impact of data-driven surveillance technologies is not borne equally by all; rather, the brunt of these technologies is disproportionately borne by already marginalized people \citep{marwick2018privacy, Eubanks2018, Roberts2019digitizing}. Although not a panacea \citep{Sloan2020participation}, participatory approaches can help ensure that the design of anti-surveillance technologies is led by those who are disproportionately targeted by surveillance. 

\bibliographystyle{abbrvnat}
\bibliography{references}

\end{document}